A bag of marbles contains 6 red marbles and 4 blue marbles. If two are selected
simultaneously, what is the probability that at least 1 blue marble is drawn?

Answer by stanbon(75887)

A bag of marbles contains 6 red marbles and 4 blue marbles. If two are selected
simultaneously, what is the probability that at least 1 blue marble is drawn?
-------------------
P(at least one is blue) = 1 - P(none are blue)
= 1 - [6C2/10C2]
= 1 - 15/45
= 1 - (1/3)
= 2/3
=================
Cheers,
Stan H.